OpenStudy (anonymous):

determine wether the graphs of y=3x+6 and 2y+6x=4 are perpendicular

OpenStudy (anonymous):

2y+6x=4 can be changed to 2y=-6x+4 and then to y=-3x+2. The graphs are not perpendicular because although the signs of the slopes are different, they are not reciprocals. The slope of of 2y+6x=4 would have to be -1/3 for the lines to be perpendicular.